OpenStudy (anonymous):

Mark bought 114 meters of fencing to enclose an area of pasture for his sheep. He plans to use all of the fencing to make a rectangle that is twice as long as it is wide. Use the information you have to create a system of equations. Let l = length and w = width. Find the length (in meters) of the area he plans to fence. @texaschic101

OpenStudy (texaschic101):

P = 2L + 2W L = 2W now sub 2W in for L in the 1st equation 114 = 2(2W) + 2W 114 = 4W + 2W 114 = 6W 114/6 = W 19 = W L = 2W L = 2(19) L = 38 check.. P = 2L + 2W 114 = 2(38) + 2(19) 114 = 76 + 38 114 = 114 (correct) Length in meters is : 38 meters

OpenStudy (anonymous):

You are correct @texaschic101 Thanks! :)

OpenStudy (texaschic101):

anytime Skiller :)

OpenStudy (anonymous):

Solve the following for w:\[2(w+2w)=114\]w = 19 2w = 38

OpenStudy (anonymous):

(w+2w) is the half perimeter. Thank you for the medal.